A coin was flipped 60 times and came up heads 38 times. (a) At the .10 level of significance, is the coin biased toward heads? Show your decision rule and calculations.
(b) Calculate a p-value and interpret it.

Answer by stanbon(57246)

A coin was flipped 60 times and came up heads 38 times.
(a) At the .10 level of significance, is the coin biased toward heads? Show your decision rule and calculations.
---
Ho: p = 1/2
Ha: p > 1/2
----------------
p-hat = 38/60 = 0.63333...
---
Critical value for one-tail z-test with alpha=10% is z=1.28
Reject Ho if test statistic is greater than 1.28
-------------------
Test statistic = z(0.63333) = (0.633333-0.5)/sqrt[(1/2)(1/2)/(60)] = 2.0655
----------
(b) Calculate a p-value and interpret it.
p-value = P(z>2.0655) = 0.01944
Interpretation: Less than 1.944% of test results could provide stronger
evidence for rejecting Ho.
======================================
Cheers,
Stan H.